# SecCodePLT: A Unified Platform for Evaluating the Security of Code GenAI

### v1

@@ Abstract
Existing works have established multiple benchmarks to highlight the security risks associated with Code GenAI. These risks are primarily reflected in two areas: a model’s potential to generate insecure code (insecure coding) and its utility in cyberattacks (cyberattack helpfulness). While these benchmarks have made significant strides, there remain opportunities for further improvement. For instance, many current benchmarks tend to focus more on a model’s ability to provide attack suggestions rather than its capacity to generate executable attacks. Additionally, most benchmarks rely heavily on static evaluation metrics (e.g., LLM judgment), which may not be as precise as dynamic metrics such as passing test cases. Furthermore, some large-scale benchmarks, while efficiently generated through automated methods, could benefit from more expert verification to ensure data quality and relevance to security scenarios. Conversely, expert-verified benchmarks, while offering high-quality data, often operate at a smaller scale. To address these gaps, we develop SecCodePLT, a unified and comprehensive evaluation platform for code GenAIs’ risks. For insecure code, we introduce a new methodology for data creation that combines experts with automatic generation. Our methodology ensures the data quality while enabling large-scale generation. We also associate samples with test cases to conduct code-related dynamic evaluation. For cyberattack helpfulness, we set up a real environment and construct samples to prompt a model to generate actual attacks, along with dynamic metrics in our environment. We conduct extensive experiments and show that SecCodePLT outperforms the state-of-the-art (SOTA) benchmark CyberSecEval in security relevance. Furthermore, it better identifies the security risks of SOTA models in insecure coding and cyberattack helpfulness. Finally, we apply SecCodePLT to the SOTA code agent, Cursor, and, for the first time, identify non-trivial security risks in this advanced coding agent. [^1] We provide data in https://huggingface.co/datasets/Virtue-AI-HUB/SecCodePLT.
@@ 1 Introduction
Code GenAI, including specific code generation models and general large language models, have shown remarkable capabilities in code generation (Chen et al., 2021; Dong et al., 2023; Austin et al., 2021; DeepSeek, 2022; Hui et al., 2024), reasoning (Gu et al., 2024), and debugging (Tian et al., 2024). Together with these exciting new capabilities comes concern over these models’ security risks. Recent research (Pearce et al., 2022; Bhatt et al., 2023) showed that code GenAI can produce insecure code, which significantly hinders the real-world deployment of AI-generated code. Moreover, these models can also be weaponized to facilitate cyberattacks.
@@
Eval. platform IC ❶ CH ❶ Metric ❷ Expert verified ❸ Risk categories
@@
[…]
To understand the security risks of code GenAI, existing works developed several benchmarks to evaluate a code generation model’s risk in producing insecure or vulnerable code (insecure coding) (Bhatt et al., 2023; 2024), as well as its potential to facilitate cyberattacks (cyberattack helpfulness) (Bhatt et al., 2024; Yuan et al., 2024). However, as demonstrated in Table 1, these benchmarks are foundationally limited. ❶ None of the existing benchmarks provide a comprehensive evaluation of both insecure coding and cyberattack helpfulness. Specifically, some early benchmarks (Pearce et al., 2022; Siddiq & Santos, 2022; Fan et al., 2020; Tihanyi et al., 2023) include only code completion tasks without instruction generation. Existing benchmarks on cyberattack helpfulness (Tihanyi et al., 2024; Liu et al., 2024c; Shao et al., 2024; Zhang et al., 2024; Bhatt et al., 2023; 2024; Wan et al., 2024) mainly prompt the model to provide attack suggestions or answer attack-related questions in natural language. Without requiring the model to generate the attack end-to-end, these benchmarks cannot assess a model’s real potential for being weaponized. ❷ All existing benchmarks leverage static-based metrics (rules (Pearce et al., 2022; Siddiq & Santos, 2022; Bhatt et al., 2023; 2024) or LLM-judgment (Bhatt et al., 2024; Yuan et al., 2024)). These methods are less precise than dynamic testing and are prone to false positives and false negatives (Thakur et al., 2024; Charoenwet et al., 2024). ❸ Existing benchmarks either lack scalability or quality. In particular, some benchmarks (Pearce et al., 2022; Siddiq & Santos, 2022) rely on manual efforts for dataset creation, which are of high quality but not scalable. Others (Bhatt et al., 2023; 2024) employ automated data creation, resulting in low-quality data that is unrelated to security-critical scenarios
To address these limitations, we introduce SecCodePLT, a novel evaluation platform designed to evaluate code GenAI in insecure coding and cyberattack helpfulness.
Insecure coding. We introduce a two-stage data creation pipeline, which enables scalability and ensures data quality (❸). Our method starts with generating a few seed samples for each selected type of vulnerability, i.e., one MITRE’s Common Weakness Enumeration (CWE) (MITRE, 2024b), and then employs LLM-based mutators to generate more data from these seeds. We include a validation step to filter out incorrect data, balancing correctness and scalability. More specifically, our seed generation begins by analyzing the vulnerability and context of a given CWE to manually create several security-related coding tasks. For each task, we generate both vulnerable and patched code versions, along with functionality and security test cases. In cases that cannot be evaluated with standard test cases, we establish detailed rules for vulnerability detection. Each task, including its description, code and tests, forms a seed. Given that all seeds are created and verified by human experts, they are guaranteed to be security-related. As detailed in Section 3, our automatic and validate process will also ensure the security relevance and correctness of newly generated data. Additionally, our samples contain both text descriptions and example code, enabling both instruction generation and code completion tasks (❶). After generating the benchmark, we further design hybrid evaluation metrics that combine dynamic testing with rule-based detection for identifying insecure code. As discussed in Chhabra & Gupta (2010), hybrid metrics are more precise than pure static metrics (❷).
Cyberattack helpfulness. We then construct a cyberattack helpfulness benchmark to evaluate a model’s capability in facilitating end-to-end cyberattacks. Following MITRE ATT&CK (MITRE, 2024a), we break down a typical cyberattack into multiple steps, treating each as a category to guide the model’s ability to perform specific aspects of an attack, such as writing exploits and deploying attacks (❸). We then design tailored prompts for each category to guide the model to generate executable attacks (❶). Finally, we create an environment with metrics to dynamically evaluate a model’s outputted attack for each category (❷).
Through extensive experiments, we first show that SecCodePLT significantly outperforms CyberSecEval in security relevance and prompt faithfulness, highlighting the quality of our benchmark. Second, we apply SecCodePLT and CyberSecEval to four SOTA open and closed-source models, showing that SecCodePLT can better reveal a model’s risk in generating insecure code. We also find that larger models tend to be less risky. Third, we evaluate the selected models using our cyberattack helpfulness benchmark and find that some widely used models can generate end-to-end attacks, indicating they can be immediately weaponized for attack and highlighting the need for better safety alignment. Finally, using our platform, we identify security risks in the SOTA code agent, Cursor (Cursor, 2024). We find that the cursor fails entirely on some critical CWEs representing code injection, access control, and data leakage prevention. To the best of our knowledge, this is the first platform to enable precise security risks evaluation and end-to-end cyberattack helpfulness assessment of code GenAI. Additionally, we are the first to reveal the security risks in Cursor.
@@ 2 Related Works
There have been several efforts to create benchmarks for training and evaluating models in code generation across various domains, such as general code generation (Chen et al., 2021; Dong et al., 2023; Austin et al., 2021), solving LeetCode challenges (DeepSeek, 2022), addressing data science problems (Lai et al., 2023), and program-based mathematical reasoning (Cobbe et al., 2021; Hendrycks et al., 2021). These datasets focus on either instruction generation or code completion. Two metrics are widely used across these datasets – CodeBLEU (Ren et al., 2020), which measures the similarity of the generated code and ground truth, and pass@K (Liu et al., 2024b), which evaluates whether the generated code can pass the predefined test cases. Beyond code generation, another line of research focuses on developing specialized datasets to enhance models’ capabilities in understanding (Gu et al., 2024; Liu et al., 2024a), debugging (Tian et al., 2024), or refining input programs (Ding et al., 2024b; Zheng et al., 2024; Yang et al., 2024; Jimenez et al., 2023).
There are also several benchmarks for testing the security of code GenAI. Most of them evaluate a model’s capability of identifying vulnerabilities (Fan et al., 2020; Ding et al., 2024a; Ullah et al., 2024). These datasets collect code snippets extracted from GitHub commits, corresponding to known CWEs or CVEs. However, most datasets have a high false positive rate, as extracting isolated code chunks without context does not always guarantee that the extracted code is truly vulnerable. CyberSecEval (Bhatt et al., 2023; 2024; Wan et al., 2024) constructs both insecure coding and cyberattack helpfulness benchmarks. To construct the insecure coding dataset, they first implement a rule-based ICD and use it to scan and extract insecure code chunks across various open-source projects. For each extracted chunk, leverage an LLM to generate a corresponding prompt that describes the code. This prompt is designed to instruct a code generation model to rewrite the code chuck. If a target model produces the original or similar insecure code, the paper deems the model as a high-risk one. This dataset has two limitations. First, the ICD often introduces false positives. Even when the ICD correctly identifies insecure code, similar to the vulnerability detection datasets discussed earlier, extracting code chunks without proper context frequently leads to false positives. As a result, a significant portion of the extracted code is not security-related or not even not even vulnerable. Second, using LLMs to generate prompts from the code does not always yield faithful descriptions. In some cases, the generated prompts are inaccurate, leading to situations where code generation models produce irrelevant code due to inaccurate descriptions. Recent research also evaluated the capability of LLMs to assist cyberattacks. For example, works like Tihanyi et al. (2024); Liu et al. (2024c) assess LLMs’ knowledge in cybersecurity through question-answering. Shao et al. (2024); Zhang et al. (2024) investigates whether LLMs can help solve offensive Capture The Flag (CTF) challenges. CyberSecEval (Bhatt et al., 2023; 2024; Wan et al., 2024) explore whether LLMs can provide suggestions for cyberattacks.
Finally, there is a line of works that assesses the risks for general LLMs (Wang et al., 2023; Sun et al., 2024), or adversarial attacks against code generation models (Wu et al., 2023; Hajipour et al., 2023; Pearce et al., 2022), which is out of our scope.
3 Key Techniques

### v2

# SeCodePLT: A Unified Benchmark for Evaluating the Security Risks and Capabilities of Code Agents

@@ Abstract
Existing benchmarks for evaluating the security risks and capabilities (e.g., vulnerability detection) of code-generating large language models (LLMs) face several key limitations: (1) limited coverage of risk and capabilities; (2) reliance on static evaluation metrics such as LLM judgments or rule-based detection, which lack the precision of dynamic analysis; and (3) a trade-off between data quality and benchmark scale. To address these challenges, we introduce a general and scalable benchmark construction framework that begins with manually validated, high-quality seed examples and expands them via targeted mutations. Our approach provides a comprehensive suite of artifacts so the benchmark can support comprehensive risk assessment and security capability evaluation using dynamic metrics. By combining expert insights with automated generation, we strike a balance between manual effort, data quality, and benchmark scale. Applying this framework to Python, C/C++, and Java, we build SeCodePLT, a dataset of more than 5.9k samples spanning 44 CWE-based risk categories and three security capabilities. Compared with state-of-the-art benchmarks, SeCodePLT offers broader coverage, higher data fidelity, and substantially greater scale. We use SeCodePLT to evaluate leading code LLMs and agents, revealing their strengths and weaknesses in both generating secure code and identifying or fixing vulnerabilities. [^1] We provide our code in https://github.com/ucsb-mlsec/SecCodePLT, data in https://huggingface.co/datasets/secmlr/SecCodePLT
@@ 1 Introduction
Code GenAI, including code LLMs and agents, has shown remarkable capabilities in general coding tasks, including code generation chen2021evaluating ; dong2023codescore ; hui2024qwen , code understanding gu2024cruxeval , and self-debugging tian2024debugbench . However, recent work demonstrated the concerning security risks of these models bhatt2023purple ; bhatt2024cyberseceval ; yuan2024rjudge . Besides, it is unclear whether their general coding capabilities can be migrated to perform more specialized security tasks, e.g., identifying and analyzing security vulnerabilities.
Existing benchmarks on evaluating a coding model’s risk have the following limitations. ❶ Existing benchmarks have limited coverage over security risks and tasks. For example, some early benchmarks pearce2022asleep ; siddiq2022seceval ; fan2020ac ; tihanyi2023formai include only code completion tasks without instruction (text-to-code) generation and other security-specific tasks (e.g., vulnerability detection). Others support only vulnerability detection ding2024vulnerability ; ullah2024llms or instruction generation he2023large ; hajipour2024codelmsec ; siddiq2022securityeval ; peng2025cweval , without a comprehensive risk and task coverage. Furthermore, many benchmarks do not even support code generation, and they test a model’s security knowledge via text-based question answering tihanyi2024cybermetric ; liu2024cyberbench ; shao2024empirical ; bhatt2023purple ; bhatt2024cyberseceval ; wan2024cyberseceval . ❷ Most existing benchmarks leverage static-based metrics (rules pearce2022asleep ; siddiq2022seceval ; bhatt2023purple ; bhatt2024cyberseceval or LLM-judgment bhatt2024cyberseceval ; yuan2024rjudge ). These methods are less precise than dynamic testing, especially LLM judgment, which relies on LLM capabilities and prompt qualities thakur2024judgingjudgesevaluatingalignment ; charoenwet2024empiricalstudystaticanalysis . ❸ There is a trade-off between data quality and scale. In particular, some benchmarks (e.g., pearce2022asleep ; siddiq2022seceval ; he2023large ; zhang2024cybench ; cyberseceval4 ) rely on manual efforts for dataset creation, which are of high quality but not scalable. Others ( bhatt2023purple ; bhatt2024cyberseceval ; ding2024vulnerability ) employ automated data creation, resulting in low-quality data and data that are not related to security risks and tasks.
To address these limitations, we introduce SeCodePLT, a novel benchmark that comprehensively evaluates code GenAI’s security risks and capabilities across multiple programming languages. Technically speaking, we introduce a two-stage data creation pipeline, which balances the manual effort, data quality, and scalability (❸). Our method starts with collecting and validating high-quality seed samples for each selected type of vulnerability/risk, i.e., Common Weakness Enumeration (CWE) (mitre_cwe_2024, ), and then employs LLM-based mutators to generate more data from these seeds. During seed collection, we manually analyze a target CWE and create security-related coding scenarios. We include a comprehensive set of artifacts for each scenario, including both vulnerable and patched code, along with functionality and security test cases. These artifacts are either manually created or extracted from real-world code cases with human validation. Then, we design an automated generation and validation method that leverages LLMs to mutate the seeds to scale up our benchmark and dynamic testing to filter out incorrect data. Given that SeCodePLT includes a comprehensive set of artifacts, it supports multiple capability evaluations, including secure code generation, vulnerability detection, and patch generation. Besides, the proposed data creation pipeline is generalizable across different languages and risks (❶). Our benchmark also supports dynamic testing, which is more precise than pure rule-based or LLM judgment (❷).
@@
Benchmark Task and risk coverage ❶ Metric ❷ Languages Verified & Scales ❸
SC VD Patch # Risk Verified # Data
@@
AutoPatchBench (cyberseceval4, ) $Circle \CIRCLE$ 6 Dynamic C/C++ 136
SeCodePLT (Ours) $CIRCLE \CIRCLE$ 44 Static + Dynamic Python, C/C++, Java 1.6k (5.9k)
We apply the proposed data creation pipeline to four popular programming languages: Python, C, C++, Java, and cover in total 44 risks (CWEs). In Table 1, we show the advantage of SeCodePLT over SOTA representative benchmarks in risk and capability coverage, data quality, and metrics. We select the earliest benchmark AsleepAtTheKeyboard pearce2022asleep , two representative benchmark in vulnerability detection: PrimeVul ding2024vulnerability , SecLLMHolmes ullah2024llms , four representative secure code generation benchmarks: SVEN he2023large , CyberSecEval bhatt2023purple , CodeLMSec hajipour2024codelmsec and BaxBench vero2025baxbench , and a patching dataset AutoPatchBench cyberseceval4 . We do not include SWE-bench jimenez2023swe , as it is not security-specific. Then, we conduct experiments to validate the data quality of SeCodePLT in security relevance, prompt faithfulness, and testing case coverage. Finally, we apply SeCodePLT to evaluate five SOTA open and closed-source models and the Cursor agent in three capabilities: secure code generation, vulnerability detection, and patch generation. Our experiment reveals the security risks of these models and agents during code generation, including code injection, access control, data leakage prevention, etc. We further show that these models are limited in identifying vulnerable code and generating security patches, highlighting the limitations of general-purpose LLMs in security-specific capabilities. In Appendix G, we further build a simulated network environment and demonstrate that SOTA code GenAI can be weaponized to generate end-to-end cyber attacks.
Contributions. We propose a novel benchmark construct pipeline for code LLM that balances scalability and data quality. We construct the first large-scale benchmark (5.9k) that enable comprehensive security risks evaluation of code GenAI using dynamic metrics across multiple programming languages. We apply SeCodePLT to multiple SOTA LLMs and uncover their limitations in multiple security capabilities.
@@ 2 Related Works
Most existing code-related benchmarks are about general code generation capabilities  dong2023codescore ; austin2021program ; zhuo2024bigcodebench ; leetcode ; lai2023ds (e.g., solving LeetCode challenges leetcode and addressing data science problems lai2023ds ), as well as code understanding gu2024cruxeval ; liu2024codemind and self-debugging tian2024debugbench ; ding2024cycle ; yang2024intercode ; jimenez2023swe . The main security-related benchmarks are two-fold: secure coding, which evaluates a model’s security awareness during code generation, and vulnerability detection, which evaluates a model’s capabilities in identifying vulnerable code.
Secure coding. These benchmarks construct instruction generation or code completion tasks for LLMs from vulnerable code of known CWEs. They then test LLMs with these tasks and evaluate if the generated code is vulnerable, which can reveal the model’s awareness of security risks during code generation. SOTA benchmarks include CyberSecEval bhatt2023purple ; bhatt2024cyberseceval ; wan2024cyberseceval , CodeLMSec hajipour2024codelmsec , LLMSecEval tony2023llmseceval , and RMCBench chen2024rmcbench . As shown in Table 1, CyberSecEval and CodeLMSec do not support dynamic testing and have limited coverage of risks and capabilities. Besides, without human validation, their coding tasks are not all related to security. For example, CyberSecEval uses a rule-based insecure coding detector (ICD) to extract insecure code chunks and leverages an LLM to generate prompts that describe the chunks. Here, the ICD introduces false positives. Even when the ICD is correct, extracting code chunks without proper context (background of the entire codebase and related functions) frequently leads to false positives. Compared to LLMSecEval and RMCBench, SeCodePLT covers more CWEs, provides structured inputs, ensures security relevance through manual verification, and includes test cases for dynamic evaluation. There are some early works on secure code generation through prompting, training, or decoding techniques he2023large ; he2024instruction ; fu2024constrained . They also construct some datasets (e.g., SVEN he2023large ) to evaluate their techniques on relatively small models. These datasets share similar limitations to the ones discussed above.
Vulnerability detection. Regarding the security capability, most existing benchmarks focus on vulnerability detection fan2020ac ; ding2024vulnerability ; ullah2024llms . Among these benchmarks, PrimeVul ding2024vulnerability , which aggregates several existing datasets, is currently the largest and most comprehensive benchmark. To achieve a large scale, PrimeVul employs a fully automated approach for vulnerable and benign code chunks extraction. As discussed above, isolated code chunks without the necessary context will introduce ambiguity for vulnerability. For example, a function may accept an argument whose definition is missing. In this case, even if the function is benign, the model may still be conservative and flag it as potentially vulnerable, as the necessary information about the argument is missing. SecLLMHolmes ullah2024llms , on the other hand, constructs a high-quality dataset with human validation. However, the dataset is very small. SeCodePLT leverages a semi-automated approach for data creation, which balances the trade-off. Note that there are more benchmarks on non-LLM vulnerability detection and vulnerability reproduction (VulHub (vulhub, ), HackTheBox Hackthebox , OWASP owasp ), which is not our focus.
Other security-related benchmarks. Recent research also evaluated the capability of LLMs to assist cyberattacks. For example, CyberMetric tihanyi2024cybermetric , CyberBench liu2024cyberbench , and CyberSecEval (bhatt2023purple, ; bhatt2024cyberseceval, ; wan2024cyberseceval, ) assess LLMs’ knowledge in cyberattacks through question-answering. RedCode guo2024redcode and RMCBench chen2024rmcbench evaluate the model in generating malware. Existing works also construct cyber ranges, such as MITRE’s Caldera caldera and IBM Cyber Range xforce , which are mainly designed to interact with humans rather than LLMs. Although it is not the main focus of this paper, we construct a simulated network system and tasks for evaluating LLMs in generating end-to-end cyber attacks. Different from existing text-based benchmarks, our design enables actual attack generation and evaluation (See Appendix G). Besides cyberattacks, there are also CTF benchmarks (Cybench zhang2024cybench and NYU CTF benchmark shao2024empirical ), penetration test benchmarks (gioacchini2024autopenbench, ), and benchmarks for backend vero2025baxbench and patching jimenez2023swe ; cyberseceval4 . These benchmarks are complementary to our benchmarks in terms of security capabilities. Note that the adversarial attacks against (code) LLMs sun2024trustllm ; wu2023deceptprompt ; pearce2022asleep are out of our scope.
3 Benchmark Construction Methodology